OpenStudy (anonymous):

Find the equation of a quadratic function f whose graph has a vertical axis of symmetry x = -2, the range of f is given by the interval [4 , +infinity) and f(2) = 8.

OpenStudy (anonymous):

you are being told in a round about way that the vertex is \((-2,4)\)

OpenStudy (anonymous):

this means it looks like \(y=a(x+2)^2+4\) the only number you do not know is \(a\)

OpenStudy (anonymous):

but you can solve for it, since \[f(2)=8\]you know \(8=a(2+2)^2+4\) and so \[16a+4=8\] \[16a=4\] \[a=\frac{1}{2}\]

OpenStudy (anonymous):

how come the vertex is (-2,4)? what does the vertical axis of symmetry x= -2 mean?

OpenStudy (anonymous):



OpenStudy (anonymous):

means a) the parabola is symmetric about the line \(x=-2\) and b) the first coordinate of the vertex is \(-2\)

OpenStudy (anonymous):

the fact that the range is \([4\infty)\) tells you the second coordinate of the vertex is \(4\)

OpenStudy (anonymous):

I see...that's what I don't know before. I just know that there is a line x=-2

OpenStudy (anonymous):

 if I may ask more, if the graph's like this, how would you say it in words?